Find the perimeter of a square whose diagonal measures 12in.
Answer by checkley75(3666)

The 2 sides of a square & the diagonal form a right triangle with equal sides.
Thus the equation for the diagonal is:
d^2=s^2+s^2
12^2=2s^2
2s^2=144
s^2=144/2
s=sqrt72
s=8.49 for the sides.
thus the perimeteris=4*side or:
4*8.49=33.96 is the perimeter.